OpenStudy (anonymous):

nico was stunned when he opened the mail.Here was a check for $332.50 this is the interest he earned on his investment.One year ago Nico had been given 5,000.His parents made him put it in the bank and save it for college,but the said he could keep the interest it earned each year.A portion of his money is deposited in a high rate account that earned 8% annual interest.The rest is in a regular account that earned 3% this year.Both accounts pay simple interest computed at the end of the year.how much of Nicos money was invested in each account.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Do you know the formula for "simple interest" ?

OpenStudy (anonymous):

I=prt

OpenStudy (phi):

now we have to figure out how to use it for your problem.

OpenStudy (anonymous):

ok

OpenStudy (phi):

How much time are we talking about? can you figure that out by reading the question ?

OpenStudy (anonymous):

year

OpenStudy (phi):

yes, 1 year. that means t= 1 in your formula (for both rates) next, we have two rates. what are they ?

OpenStudy (anonymous):

8 and 3

OpenStudy (phi):

I think we have to be more careful. not just 8 and 3. what about the percent ?

OpenStudy (phi):

Do you understand that 8% is different from just 8? And when we do these problems we have to write the rate as a *decimal* number. Can you write 8% as a decimal?

OpenStudy (anonymous):

i kno

OpenStudy (phi):

Can you write 8% as a decimal?

OpenStudy (anonymous):

0.08

OpenStudy (phi):

ok. Now let's write what we know about the interest for part of the money invested at 8%: I = P R T T=1 (1 year) R = 0.08 (8% per year) P = ? we don't know (and it's what we have to find to answer the question) let P = x (x for unknown) I = x * 0.08 * 1 multiplying by 1 does not do anything, so let's simplify to just I = 0.08x (this means 0.08 * x) ok so far ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

If we call the amount invested at 3% "y" , can you try writing the same type of equation for the amount invested at 3% ?

OpenStudy (anonymous):

the equation for 3% is 0.03x

OpenStudy (phi):

ok but it's better to say Interest= 0.03x just to help remember what it stands for. Now we use this idea: the sum of both interests add up to what number ?

OpenStudy (phi):

and we have to be more careful. x is the amount invested at 8% that amount invested at 3% is probably not the same amount, so we should call it y (otherwise we will get confused) so I= 0.8x I2 = 0.03y (I2 is to show it's a different interest) the total interest we get is I+I2 = ? (read the problem)

OpenStudy (anonymous):

0.8x+0.03x

OpenStudy (phi):

almost. do you understand that we should not use 0.03x , it should be 0.03y ?

OpenStudy (phi):

and somehow 0.08 was wrongly changed to 0.8 oops!

OpenStudy (phi):

also a check for $332.50 this is the interest he earned on his investment. in other words the total of the two interests is that amount. can you try again?

OpenStudy (anonymous):

0.08x+0.03y=332.50

OpenStudy (phi):

ok, now it looks good. next, any idea what x+y add up to ? (remember x and y are the amounts we invest)

OpenStudy (anonymous):

0.11

OpenStudy (phi):

don't confuse the interest rates. x and y are the principal (amount of money) how much money was invested? (read the problem, it tells you)

OpenStudy (anonymous):

5,000

OpenStudy (phi):

do you see x+y= 5000 and we split the 5000 into x for the 8%, and y for the 3% ?

OpenStudy (phi):

Here is what we know 0.08x+0.03y=332.50 x + y = 5000 do you know how to solve a system of equations ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@phi

OpenStudy (phi):

we can use substitution. start with x+y= 5000 can you "solve for y" ? (add -x to both sides)

OpenStudy (anonymous):